A data set has the following points:
(1,1), (2,5), (3,10)
The slope of the linear regression line is 4.5. Find the y-intercept of the linear regression.

Answer by stanbon(75887)

A data set has the following points:
(1,1), (2,5), (3,10)
The slope of the linear regression line is 4.5. Find the y-intercept of the linear regression.
----
Note: The three points are not on a single line.
You can see that by calculating the slope between
each pair of points.
============================
Use the point (1,1) and slope = 9/2 to find "b":
y = mx + b
---
1 = (9/2)1 + b
b = (2/2)-(9/2)
b = -7/2
===============
Cheers,
Stan H.
===============